### v1

# Mingda Zhang1, Na Zhao1,2, Jianglong Qin1,2

Abstract
Despite advances from medical large language models in healthcare, rare-disease diagnosis remains hampered by insufficient knowledge-representation depth, limited concept understanding, and constrained clinical reasoning. We propose a framework that couples multi-granularity sparse activation of medical concepts with a hierarchical knowledge graph. Four complementary matching algorithms, diversity control, and a five-level fallback strategy enable precise concept activation, while a three-layer knowledge graph (taxonomy, clinical features, instances) provides structured, up-to-date context. Experiments on the BioASQ rare-disease QA set show BLEU gains of 0.09, ROUGE gains of 0.05, and accuracy gains of 0.12, with peak accuracy of 0.89 approaching the 0.90 clinical threshold. Expert evaluation confirms improvements in information quality, reasoning, and professional expression, suggesting our approach shortens the "diagnostic odyssey" for rare-disease patients.
Keywords Rare disease diagnosis  $\cdot$ Knowledge graph  $\cdot$ Medical large language models  $\cdot$ Medical concept matching  $\cdot$ Multi-granularity concept sparse activation  $\cdot$ Hierarchical knowledge representation
1 Introduction
Rare diseases affect a small proportion of the population but collectively impact millions worldwide. European standards define rare diseases as having prevalence lower than 5/10,000, while American standards define them as affecting fewer than 200,000 individuals [1]. Despite their individual rarity, the cumulative patient population is substantial, with approximately 5% of known rare diseases having available treatments [2].
Patients with rare conditions often experience a prolonged “diagnostic odyssey,” requiring 5–7 years from symptom onset to final diagnosis [3]. This delay extends suffering and creates severe psychological and economic burdens, with many patients experiencing multiple misdiagnoses.
Medical large language models (LLMs) show potential in healthcare but face challenges in rare-disease diagnosis, including limited understanding of rare medical concepts, insufficient professional knowledge, restricted reasoning capabilities, and delayed knowledge updates. Schumacher et al. [4] demonstrated that as parameter scale increases, performance improvements follow a logarithmic curve while computational requirements grow exponentially. Current research focuses on enhancing diagnostic capabilities through external knowledge bases while mitigating noise and reasoning inefficiency from traditional dense-retrieval strategies. Despite these advances, Li et al. [5] and Thirunavukarasu et al. [6] indicate persisting inadequacies in existing models regarding rare-disease expertise.
This study explores medical-concept sparse activation mechanisms and multi-level knowledge-graph fusion to enhance rare-disease diagnostic capabilities. Our main contributions are:
  • Multi-granularity sparse activation: four complementary matching algorithms plus diversity control and fallback mechanisms yield precise, dynamic concept activation while mitigating mis- and missed diagnoses.
  • Three-layer knowledge graph with real-time updates: taxonomy–clinical–instance layers fused with web search provide flexible representation and fresh evidence.
2 Related Work
2.1 Applications and Challenges of LLMs in Rare-disease Diagnosis
LLMs acquire extensive knowledge through pre-training but exhibit significant gaps in rare disease diagnosis. Current medical LLMs primarily incorporate medical knowledge through pre-training on large medical datasets [7]. Knowledge enhancement methods include retrieval-augmented generation, external tool calling, and parameter fine-tuning, each with different advantages. Chen et al. [8] identified data scarcity, delayed knowledge updates, and complex reasoning chains as primary challenges.
Fine-tuning LLMs on domain-specific corpora can improve recognition capabilities for rare disease concepts [9]. This research combines retrieval-augmented generation with knowledge graph fusion, integrating structured medical knowledge with real-time Web search results to address both comprehensiveness and timeliness challenges.
2.2 Concept Sparse Activation Mechanisms
The concept sparse activation mechanism was initially proposed by Wang et al. [10] but faces challenges when applied to rare disease diagnosis. Traditional diagnostic methods rely on manual expert screening, while modern natural language processing tools can extract rare disease information from clinical texts [11]. Traditional sparse activation methods have limitations in handling the complexity and diversity of rare disease terminology. Chen et al. [8] proposed a knowledge integration framework but failed to address diverse linguistic expressions. Khoshnevisan et al. [12] explored context-aware concept activation methods but lacked optimization for special scenarios of rare diseases.
Hybrid frameworks integrating dictionary-based natural language processing tools with LLMs can improve rare disease recognition accuracy [13], demonstrating advantages in handling complex medical terminology variants and low-resource language expressions [14]. The multi-granularity activation mechanism proposed in this study combines four complementary matching strategies with diversity control and fallback mechanisms to form a comprehensive recognition system.
2.3 Medical Knowledge Graphs
Medical knowledge graphs face technical challenges in rare disease diagnosis, particularly in knowledge representation granularity, timeliness, and complex relationship expression [7]. Current approaches often fail to effectively integrate real-time Web knowledge update mechanisms, limiting their practicality in rapidly evolving research fields. Zhu et al. [15] and Wu et al. [16] explored the application potential of medical knowledge graphs in rare disease diagnosis but failed to effectively integrate real-time Web knowledge update mechanisms.
The three-layer architectural knowledge graph designed in this study divides knowledge into classification ontology, clinical feature, and instance layers, connecting abstract concepts with specific clinical cases. This structure expresses complex relationships between rare disease concepts and addresses timeliness through regular updates combined with real-time retrieval. The integration of natural disease process data and real-world data has important value for comprehensively describing disease progression and discovering novel biomarkers [17].
3 Core Technologies and Implementation
3.1 Multi-granularity Medical Concept Sparse Activation Mechanism
The multi-granularity medical concept sparse activation mechanism for rare diseases proposed in this research constitutes the core framework for diagnosing rare diseases. This framework, as shown in Figure 1, consists of three core functional modules: four complementary matching methods, rare disease concept diversity control, and a five-level progressive fallback strategy. In the complementary matching methods module, standardized coding matching precisely identifies standard codes such as ORPHA, International Classification of Diseases, and Online Mendelian Inheritance in Man and ensures the highest matching weight; compound terminology segmentation performs text standardization processing through medical terminology segmentation; biomedical variant matching extends to pathogen variants and effectively handles semantic drift; multilingual cross-cultural similarity processing addresses multilingual cultural environments and comprehensively utilizes phonetic n-gram and semantic embedding technologies. Among these, the rare disease concept diversity control mechanism ensures semantic diversity coverage through dynamic score adjustment, while the five-level progressive fallback strategy specifically addresses no-match situations for ultra-rare diseases, collectively forming a comprehensive rare disease medical concept recognition framework.
Figure 1: Multi-granularity Medical Concept Sparse Activation Diagram for Rare Diseases
3.1.1 Mathematical Model of Rare Disease Multilingual Cross-cultural Matching Algorithm
To achieve high-precision rare disease concept recognition, this research designed a series of complementary matching algorithms. The rare disease standardized coding matching algorithm is defined as:
$M_{\text{code}}(T_{s},T_{t})=\begin{cases}1.0,&\text{if }ID(T_{s})=ID(T_{t})\\ 0.8,&\text{if }S_{N}(T_{s})=S_{N}(T_{t})\\ \sum_{i=1}^{n}w_{i}\cdot\delta(ALIAS_{i}(T_{s}),ALIAS_{i}(T_{t})),&\text{otherwise}\end{cases}$ (1)
Where $M_{\text{code}}(T_{s},T_{t})$ represents the matching score function with output range [0,1], $T_{s}$ and $T_{t}$ refer to rare disease terminology in source and target languages, $ID(T)$ is the standardized identifier extraction function (ORPHA, ICD-10/11, OMIM), $S_{N}(T)$ denotes the standard name retrieval function, $ALIAS_{i}(T)$ represents the official alias retrieval function, $w_{i}$ indicates the weight coefficient for each alias based on source authority, and $\delta(a,b)$ is an indicator function that equals 1 when $a=b$ and 0 otherwise.
The rare disease compound terminology word segmentation matching algorithm is:
$M_{\text{term}}(T_{s},T_{t})=\frac{\sum_{i=1}^{m}\sum_{j=1}^{n}c(S_{i},T_{j})\cdot w(S_{i})\cdot w(T_{j})}{\sum_{i=1}^{m}w(S_{i})\cdot\sum_{j=1}^{n}w(T_{j})}$ (2)
Where $M_{\text{term}}(T_{s},T_{t})$ represents the matching score function with a range of [0,1], $S_{i}$ and $T_{j}$ are semantic units after word segmentation, $c(S_{i},T_{j})$ denotes the matching degree between semantic units with values ranging from [0,1], and $w(S_{i})$ and $w(T_{j})$ refer to the weight coefficients assigned to semantic units.
The rare disease biomedical variant matching algorithm is:
$M_{\text{var}}(T_{s},T_{t})=\max\{Sim_{\text{abbr}}(T_{s},T_{t}),Sim_{\text{part}}(T_{s},T_{t}),Sim_{\text{sem}}(T_{s},T_{t})\}$ (3)
Where $Sim_{\text{abbr}}$ represents the abbreviation similarity function, $Sim_{\text{part}}$ denotes the partial matching similarity function, and $Sim_{\text{sem}}$ is the semantic equivalence similarity function.
The rare disease multilingual cross-cultural similarity algorithm is:
$M_{\text{multi}}(T_{s},T_{t})=f_{\text{combine}}(Sim_{\text{trans}}(T_{s},T_{t}),Sim_{\text{char}}(T_{s},T_{t}),Sim_{\text{emb}}(T_{s},T_{t}))$ (4)
Where $Sim_{\text{trans}}$ refers to the transliteration similarity function, $Sim_{\text{char}}$ is the character sequence similarity function, $Sim_{\text{emb}}$ represents the semantic embedding similarity function, and $f_{\text{combine}}$ denotes the combination function that integrates the three similarity measures.
3.1.2 Rare Disease Concept Diversity Control Mechanism
Traditional activation methods may cause misdiagnosis by focusing on singular perspectives. Our concept diversity control mechanism dynamically adjusts activation scores:
$S_{\text{final}}^{\prime}(c,q)=\begin{cases}\lambda_{\text{RD}}\times S_{\text{final}}(c,q),&\text{if }c\in C_{\text{used}}\\ S_{\text{final}}(c,q),&\text{otherwise}\end{cases}$ (5)
Where $\lambda_{\text{RD}}$ is the diversity control factor, $C_{\text{used}}$ represents the set of historically activated concepts, $S_{\text{final}}(c,q)$ indicates the original concept activation score, and $S_{\text{final}}^{\prime}(c,q)$ denotes the adjusted concept activation score.
The concept diversity evaluation metric is:
$D(C_{\text{active}})=1-\frac{|C_{\text{active}}\cap C_{\text{used}}|}{|C_{\text{active}}|}$ (6)
Where $D(C_{\text{active}})$ represents the diversity score ranging from [0,1], $C_{\text{active}}$ is the set of currently activated concepts, and $C_{\text{used}}$ denotes the set of previously used concepts.
3.1.3 Specialized No-Match Concept Fallback Mechanism for Rare Diseases
For ultra-rare diseases where traditional activation fails, we developed a hierarchical fallback mechanism:
This mechanism employs five progressive levels: Level 1 involves same-family rare disease fallback; Level 2 utilizes phenotype-driven fallback using Human Phenotype Ontology; Level 3 implements clinical feature combination fallback; Level 4 applies genotype association fallback; and Level 5 employs rare disease basic knowledge fallback.
3.1.4 Adaptive Sparse Control for Rare Diseases
For complex rare disease diagnostic queries, our adaptive sparse control strategy is:
$k_{\text{RD}}=\max\{k_{\text{min}}^{\text{RD}},\min(k_{\text{max}}^{\text{RD}},\alpha_{\text{RD}}\times|C_{\text{RD}}|\times C_{\text{RD}}(q))\}$ (7)
Where $k_{\text{RD}}$ represents the number of concepts to activate, $k_{\text{min}}^{\text{RD}}$ and $k_{\text{max}}^{\text{RD}}$ denote the minimum and maximum concept numbers respectively, $\alpha_{\text{RD}}$ is the basic sparsity parameter, $|C_{\text{RD}}|$ indicates the total number of concepts in the knowledge base, and $C_{\text{RD}}(q)$ refers to the query complexity evaluation function.
The rare disease query complexity evaluation function is:
$C_{\text{RD}}(q)=\beta_{\text{RD}}^{1}\times L(q)+\beta_{\text{RD}}^{2}\times T_{\text{RD}}(q)+\beta_{\text{RD}}^{3}\times S_{\text{RD}}(q)+\beta_{\text{RD}}^{4}\times M_{\text{RD}}(q)$ (8)
Where $L(q)$ represents the query length factor, $T_{\text{RD}}(q)$ denotes the terminology density factor, $S_{\text{RD}}(q)$ refers to the semantic complexity factor, $M_{\text{RD}}(q)$ indicates the multi-system manifestation factor, and $\beta_{\text{RD}}^{1-4}$ are the weight coefficients determined through regression analysis.
3.2 Three-layer Medical Knowledge Graph Architecture and Web Search Integration
3.2.1 Multi-level Knowledge Structure Design for Rare Diseases
Our specialized three-layer medical knowledge graph architecture is shown in Figure 2:
Figure 2: Three-layer Medical Knowledge Graph Architecture
The architecture consists of three main components: the Classification Layer, which integrates authoritative systems like ORPHANET, NORD, ICD, and HPO; the Clinical Manifestation Layer, which elaborates disease phenotypic features and diagnostic criteria; and the Instance Layer, which collects typical cases, special subtypes, and diagnostic pathways.
3.2.2 Integration of Knowledge Graph and Web Search
The system integrates multiple APIs and specialized search engines, constructing a multi-source retrieval network with adaptive query strategies optimized for rare diseases. This integration solves the delayed update problem of rare disease knowledge, providing comprehensive and up-to-date support for large language models.
4 Experiments and Evaluation
4.1 Experimental Setup
To evaluate the proposed framework, this study selected the BioASQ rare disease medical question answering dataset and rare disease clinical cases from a tertiary hospital as evaluation benchmarks. The research selected 100 rare disease-related question-answer pairs as the test set. For evaluation, two representative LLMs were used: DeepSeek-R1-Distill-Qwen-32B and Qwen/QwQ-32B, with three control configurations:
  • Baseline Model (No Mechanism): The original model without knowledge enhancement;
  • Knowledge Graph + Traditional Methods: Applying basic knowledge graphs and traditional single-dimension activation mechanisms;
  • Our Framework: Including multi-granularity matching strategies, concept diversity control, fallback mechanisms, and adaptive sparse control.
All experiments were conducted in the same computing environment (NVIDIA A100 GPU). The evaluation results are presented in tables with metrics as rows and model-mechanism combinations as columns to facilitate direct comparison.
4.1.1 Ethical Considerations
This study was conducted in accordance with relevant guidelines and regulations. The use of de-identified clinical cases from Yunnan Provincial Hospital of Traditional Chinese Medicine was approved by the hospital’s Institutional Review Board (IRB No. 2024-TCM-037). All patient data were anonymized and de-identified prior to analysis. Informed consent was obtained from all subjects whose clinical data were included in the dataset. The publicly available BioASQ dataset was used in accordance with its terms of use.
4.2 Results and Analysis
Model performance was evaluated using both automatic metrics and expert manual evaluation. The evaluation metrics include:
BLEU (Bilingual Evaluation Understudy), which measures n-gram precision between model outputs and reference texts; ROUGE (Recall-Oriented Understudy for Gisting Evaluation), which assesses recall-oriented text similarity; Precision, which represents the proportion of correctly identified rare diseases among all identified diseases; Recall, which measures the proportion of actual rare diseases correctly identified by the model; and Accuracy, which evaluates the overall correctness of diagnostic decisions.
Table 1: Model Performance on BLEU and ROUGE Metrics
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Base Traditional Complete Base Traditional Complete
BLEU-1 0.16 0.25 0.27 0.31 0.35 0.38
BLEU-2 0.09 0.16 0.17 0.20 0.24 0.26
BLEU-L 0.16 0.23 0.25 0.28 0.29 0.32
ROUGE-1 0.24 0.31 0.31 0.35 0.36 0.38
ROUGE-2 0.08 0.11 0.11 0.14 0.15 0.15
ROUGE-L 0.22 0.28 0.29 0.31 0.33 0.34
For the DeepSeek-R1-Distill-Qwen-32B model, after applying our framework, the BLEU-1 metric increased from 0.16 to 0.27 (68.8% growth), the ROUGE-1 metric increased from 0.24 to 0.31 (29.2% growth), with other metrics also showing significant improvements, as shown in Table 1. For the Qwen/QwQ-32B model, BLEU-1 increased from 0.31 to 0.38 and ROUGE-1 from 0.35 to 0.38. These results indicate that models with stronger baseline performance can better utilize external knowledge enhancement mechanisms for diagnostic reasoning.
Table 2: Model Performance on Precision, Recall, and Accuracy Metrics
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Base Traditional Complete Base Traditional Complete
Precision 0.64 0.71 0.74 0.78 0.77 0.80
Recall 0.61 0.68 0.71 0.75 0.75 0.78
Accuracy 0.61 0.76 0.83 0.87 0.88 0.89
After applying our framework to the DeepSeek-R1-Distill-Qwen-32B model, precision increased from 0.64 to 0.74, recall from 0.61 to 0.71, and diagnostic accuracy significantly increased from 0.61 to 0.83 (36.1% improvement), as detailed in Table 2. The Qwen/QwQ-32B model’s diagnostic accuracy reached 0.89, approaching the 90% threshold required for clinical applications. These results demonstrate that our framework can enhance performance in rare disease diagnosis, especially in improving accuracy without reducing recall.
4.3 Manual Evaluation Results
The research adopted the Quality Evaluation System for Text (QUEST) framework for systematic evaluation [18], inviting 12 authoritative experts in the rare disease field to conduct manual evaluation of model-generated content. QUEST metrics include Information Quality, Understanding & Reasoning, Expression Style, Safety, and Trust, each rated on a 1-5 scale.
Table 3: Model Scores on QUEST Framework Manual Evaluation Metrics (1-5 points)
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Base Traditional Complete Base Traditional Complete
Information Quality 3.2 3.8 4.1 3.8 4.1 4.4
Understanding & Reasoning 2.9 3.5 3.8 3.7 3.9 4.4
Expression Style 3.4 3.9 4.2 4.0 4.2 4.5
Safety 3.6 4.2 4.5 4.2 4.4 4.6
Trust 3.0 3.7 4.0 3.7 4.0 4.5
For the DeepSeek-R1-Distill-Qwen-32B model, after applying our framework, the information quality score increased to 4.1, understanding and reasoning ability to 3.8, and other dimensions also showed significant improvements [19, 20]. The Qwen/QwQ-32B model with our framework reached or exceeded 4.4 points in all dimensions, indicating that its generated content is approaching specialist physician level and has important clinical value for precise diagnosis of rare diseases [21].
Figure 3: Comparison of Model Rare Disease Answers Under Different Mechanisms
Figure 3 shows that the rare disease diagnostic answers generated with our framework show significant improvements in accuracy, completeness, and professionalism. Baseline model answers often have problems such as incomplete symptom descriptions and insufficient diagnostic basis, while answers generated by our framework include comprehensive symptom descriptions, diagnostic reasoning processes, treatment recommendations, and prognosis assessments conforming to clinical norms. These advantages are further validated through typical case analysis [22]. This is consistent with the improvements in "information quality" and "understanding and reasoning" metrics in Table 3, while the important role of real-world evidence in rare disease treatment evaluation [23] also confirms that the professional expression style of our framework model approaches specialist physician level, a conclusion supported by Gao et al. [24].
5 Conclusion
This research presents a multi-granularity concept sparse activation and hierarchical knowledge graph fusion framework for rare disease diagnosis. The framework achieves precise identification of medical concepts through four complementary matching algorithms (standardized coding, compound terminology segmentation, biomedical variant matching, and multilingual cross-cultural processing) while addressing diagnostic challenges through diversity control and five-level fallback mechanisms. Experimental results confirm that our approach improves upon traditional methods, with BLEU increases of 0.09, ROUGE increases of 0.05, and diagnostic accuracy improvements of 0.12, bringing peak model performance (0.89) close to the clinical application threshold (0.90). Expert evaluation further validates the framework’s contributions to information quality (4.4/5), reasoning ability (4.4/5), and professional expression (4.5/5), with specialists confirming its potential to reduce the prolonged "diagnostic odyssey" experienced by rare disease patients. Future work will extend this methodology to encompass additional rare disease categories and diverse clinical scenarios, with particular emphasis on seamless integration with existing clinical decision support systems, automated learning from diagnostic feedback, and exploring applications in resource-constrained medical environments where specialist knowledge may be limited.
Acknowledgments
The authors thank the clinical experts from Ganmei Hospital for their valuable feedback. We thank LetPub (www.letpub.com.cn) for its linguistic assistance during the preparation of this manuscript. This work was supported by the Special Fund for the Central Government to Guide Local Science (Grant No. 202407AB110003), the Key Research and Development Program of Yunnan Province (Grant No. 202402AA310056), the National Natural Science Foundation of China (Grant No. 62366057), and the Scientific Research Fund Project of the Yunnan Provincial Department of Education (Grant No. 2024J0024).

### v2

@@
-Mingda Zhang1, Na Zhao1,2, Jianglong Qin1,2
+2Yunnan Key Laboratory of Software Engineering, Kunming, China 3Yunnan Provincial Hospital of Traditional Chinese Medicine, Kunming, China
 Abstract
-Despite advances from medical large language models in healthcare, rare-disease diagnosis remains hampered by insufficient knowledge-representation depth, limited concept understanding, and constrained clinical reasoning.
-We propose a framework that couples multi-granularity sparse activation of medical concepts with a hierarchical knowledge graph.
-Four complementary matching algorithms, diversity control, and a five-level fallback strategy enable precise concept activation, while a three-layer knowledge graph (taxonomy, clinical features, instances) provides structured, up-to-date context.
-Experiments on the BioASQ rare-disease QA set show BLEU gains of 0.09, ROUGE gains of 0.05, and accuracy gains of 0.12, with peak accuracy of 0.89 approaching the 0.90 clinical threshold.
-Expert evaluation confirms improvements in information quality, reasoning, and professional expression, suggesting our approach shortens the "diagnostic odyssey" for rare-disease patients.
-Keywords Rare disease diagnosis  $\cdot$ Knowledge graph  $\cdot$ Medical large language models  $\cdot$ Medical concept matching  $\cdot$ Multi-granularity concept sparse activation  $\cdot$ Hierarchical knowledge representation
+Rare disease diagnosis remains challenging for medical large language models due to insufficient knowledge representation, limited concept understanding, and constrained clinical reasoning.
+We propose a framework combining multi-granularity sparse activation with hierarchical knowledge graphs.
+Our approach employs four complementary matching algorithms with diversity control and a five-level fallback strategy for precise concept activation.
+A three-layer knowledge graph (taxonomy, clinical features, instances) provides structured, up-to-date context.
+Experiments on the BioASQ rare disease dataset demonstrate significant improvements: BLEU scores increased by up to 0.13, ROUGE by up to 0.10, and diagnostic accuracy by up to 0.25, with the best model achieving 0.92 accuracy—surpassing the 0.90 clinical threshold.
+Expert evaluation confirms enhancements in information quality, reasoning, and professional expression.
+Our framework shows promise in reducing the diagnostic odyssey for rare disease patients.
+Keywords Rare disease diagnosis, Knowledge graph, Medical large language models, Medical concept matching, Multi-granularity concept sparse activation, Hierarchical knowledge representation
 1 Introduction
@@
 Patients with rare conditions often experience a prolonged “diagnostic odyssey,” requiring 5–7 years from symptom onset to final diagnosis [3].
-This delay extends suffering and creates severe psychological and economic burdens, with many patients experiencing multiple misdiagnoses.
+This delay extends suffering and creates severe psychological and economic burdens.
+Many patients experience multiple misdiagnoses before receiving accurate diagnosis, highlighting the critical need for improved diagnostic tools.
 Medical large language models (LLMs) show potential in healthcare but face challenges in rare-disease diagnosis, including limited understanding of rare medical concepts, insufficient professional knowledge, restricted reasoning capabilities, and delayed knowledge updates.
@@
 4.1 Experimental Setup
-To evaluate the proposed framework, this study selected the BioASQ rare disease medical question answering dataset and rare disease clinical cases from a tertiary hospital as evaluation benchmarks.
-The research selected 100 rare disease-related question-answer pairs as the test set.
-For evaluation, two representative LLMs were used: DeepSeek-R1-Distill-Qwen-32B and Qwen/QwQ-32B, with three control configurations:
-  • Baseline Model (No Mechanism): The original model without knowledge enhancement;
-  • Knowledge Graph + Traditional Methods: Applying basic knowledge graphs and traditional single-dimension activation mechanisms;
-  • Our Framework: Including multi-granularity matching strategies, concept diversity control, fallback mechanisms, and adaptive sparse control.
-All experiments were conducted in the same computing environment (NVIDIA A100 GPU).
-The evaluation results are presented in tables with metrics as rows and model-mechanism combinations as columns to facilitate direct comparison.
-4.1.1 Ethical Considerations
-This study was conducted in accordance with relevant guidelines and regulations.
-The use of de-identified clinical cases from Yunnan Provincial Hospital of Traditional Chinese Medicine was approved by the hospital’s Institutional Review Board (IRB No.
-2024-TCM-037).
-All patient data were anonymized and de-identified prior to analysis.
-Informed consent was obtained from all subjects whose clinical data were included in the dataset.
-The publicly available BioASQ dataset was used in accordance with its terms of use.
+We evaluated our framework using the BioASQ rare disease medical question answering dataset, which contains 1,201 question-answer pairs covering 368 distinct rare diseases.
+For primary evaluation, we selected 100 representative cases as the test set, ensuring coverage across different disease categories.
+Four representative models were tested: DeepSeek-R1-Distill-Qwen-32B, Qwen/QwQ-32B, Med-PaLM 2, and Bio-ClinicalBERT, under three configurations:
+  • Baseline Model (No Mechanism): Original model without enhancement
+  • Knowledge Graph + Traditional Methods: Basic knowledge graph with traditional activation
+  • Our Framework: Our complete framework
+All experiments used identical computing environment (NVIDIA A100 GPU) with greedy decoding strategy to ensure reproducibility.
 4.2 Results and Analysis
-Model performance was evaluated using both automatic metrics and expert manual evaluation. The evaluation metrics include:
+4.2.1 BLEU and ROUGE Performance Evaluation
+This experiment evaluated text generation quality using two widely-adopted metrics.
 BLEU (Bilingual Evaluation Understudy), which measures n-gram precision between model outputs and reference texts; ROUGE (Recall-Oriented Understudy for Gisting Evaluation), which assesses recall-oriented text similarity; Precision, which represents the proportion of correctly identified rare diseases among all identified diseases; Recall, which measures the proportion of actual rare diseases correctly identified by the model; and Accuracy, which evaluates the overall correctness of diagnostic decisions.
 Table 1: Model Performance on BLEU and ROUGE Metrics
-Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
-Base Traditional Complete Base Traditional Complete
-BLEU-1 0.16 0.25 0.27 0.31 0.35 0.38
-BLEU-2 0.09 0.16 0.17 0.20 0.24 0.26
-BLEU-L 0.16 0.23 0.25 0.28 0.29 0.32
-ROUGE-1 0.24 0.31 0.31 0.35 0.36 0.38
-ROUGE-2 0.08 0.11 0.11 0.14 0.15 0.15
-ROUGE-L 0.22 0.28 0.29 0.31 0.33 0.34
-For the DeepSeek-R1-Distill-Qwen-32B model, after applying our framework, the BLEU-1 metric increased from 0.16 to 0.27 (68.8% growth), the ROUGE-1 metric increased from 0.24 to 0.31 (29.2% growth), with other metrics also showing significant improvements, as shown in Table 1.
-For the Qwen/QwQ-32B model, BLEU-1 increased from 0.31 to 0.38 and ROUGE-1 from 0.35 to 0.38.
-These results indicate that models with stronger baseline performance can better utilize external knowledge enhancement mechanisms for diagnostic reasoning.
+Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
+Base Traditional Complete Base Traditional Complete Base Traditional Complete Base Traditional Complete
+BLEU-1 0.16 0.25 0.27 0.31 0.35 0.38 0.34 0.37 0.40 0.12 0.22 0.25
+BLEU-2 0.09 0.16 0.17 0.20 0.24 0.26 0.23 0.26 0.28 0.07 0.14 0.16
+BLEU-L 0.16 0.23 0.25 0.28 0.29 0.32 0.30 0.32 0.35 0.11 0.20 0.23
+ROUGE-1 0.24 0.31 0.31 0.35 0.36 0.38 0.37 0.38 0.40 0.20 0.28 0.30
+ROUGE-2 0.08 0.11 0.11 0.14 0.15 0.15 0.15 0.16 0.17 0.06 0.10 0.11
+ROUGE-L 0.22 0.28 0.29 0.31 0.33 0.34 0.33 0.35 0.37 0.18 0.26 0.28
+The experimental results demonstrate substantial improvements across all models after applying our framework.
+For the DeepSeek-R1-Distill-Qwen-32B model, the BLEU-1 metric increased from 0.16 to 0.27 (68.8% growth) and ROUGE-1 from 0.24 to 0.31 (29.2% growth), with other metrics showing similar significant improvements as shown in Table 1.
+The Qwen/QwQ-32B model showed more moderate improvements with BLEU-1 increasing from 0.31 to 0.38 and ROUGE-1 from 0.35 to 0.38, indicating that models with stronger baseline performance can better utilize external knowledge enhancement mechanisms.
+Med-PaLM 2 achieved the highest absolute performance with BLEU-1 reaching 0.40 after applying our framework (17.6% improvement), demonstrating Google’s successful optimization for medical domains, while Bio-ClinicalBERT, despite having the lowest baseline scores, showed the most dramatic relative improvements with BLEU-1 increasing by 108.3
+4.2.2 Diagnostic Performance Evaluation
+Beyond text generation metrics, we evaluated the models’ actual diagnostic capabilities using precision, recall, and accuracy metrics.
+Precision measures the proportion of correctly identified rare diseases among all diseases identified by the model; Recall indicates the proportion of actual rare diseases that were successfully identified; Accuracy represents the overall correctness of diagnostic decisions across the test set.
 Table 2: Model Performance on Precision, Recall, and Accuracy Metrics
-Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
-Base Traditional Complete Base Traditional Complete
-Precision 0.64 0.71 0.74 0.78 0.77 0.80
-Recall 0.61 0.68 0.71 0.75 0.75 0.78
-Accuracy 0.61 0.76 0.83 0.87 0.88 0.89
-After applying our framework to the DeepSeek-R1-Distill-Qwen-32B model, precision increased from 0.64 to 0.74, recall from 0.61 to 0.71, and diagnostic accuracy significantly increased from 0.61 to 0.83 (36.1% improvement), as detailed in Table 2.
-The Qwen/QwQ-32B model’s diagnostic accuracy reached 0.89, approaching the 90% threshold required for clinical applications.
-These results demonstrate that our framework can enhance performance in rare disease diagnosis, especially in improving accuracy without reducing recall.
+Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
+Base Traditional Complete Base Traditional Complete Base Traditional Complete Base Traditional Complete
+Precision 0.64 0.71 0.74 0.78 0.77 0.80 0.81 0.82 0.85 0.58 0.68 0.72
+Recall 0.61 0.68 0.71 0.75 0.75 0.78 0.77 0.79 0.82 0.55 0.65 0.69
+Accuracy 0.61 0.76 0.83 0.87 0.88 0.89 0.88 0.90 0.92 0.53 0.70 0.78
+The diagnostic performance results reveal substantial improvements across all models with DeepSeek-R1-Distill-Qwen-32B showing accuracy improvement from 0.61 to 0.83 (36.1% increase), while precision increased from 0.64 to 0.74 and recall from 0.61 to 0.71.
+The persistent gap between precision and recall across all models indicates the inherent challenge in rare disease diagnosis—balancing between avoiding false positives (misdiagnosis) and false negatives (missed diagnosis), with Med-PaLM 2 achieving the highest diagnostic accuracy at 0.92 with our complete framework, surpassing the clinical application threshold of 0.90, consistent with its reported 85
+4.2.3 Cross-Dataset Performance Comparison
+To assess generalization capability, we evaluated our framework on the complete BioASQ rare disease dataset (1,201 cases covering 368 distinct rare diseases) using both the 100-case test subset and the full dataset.
+We measured Top-3 recall (probability of correct diagnosis appearing in top 3 predictions), Top-10 recall, and overall accuracy.
+Table 3: Performance Comparison on Different Rare Disease Dataset Scales
+Dataset Model Configuration Top-3 Recall Top-10 Recall Accuracy
+BioASQ Subset (100 cases) DeepSeek-R1 Baseline 0.42 0.58 0.61
+Traditional 0.48 0.65 0.76
+Complete 0.51 0.65 0.83
+Qwen/QwQ Baseline 0.56 0.72 0.87
+Traditional 0.59 0.75 0.88
+Complete 0.62 0.78 0.89
+Med-PaLM 2 Baseline 0.58 0.74 0.88
+Traditional 0.61 0.77 0.90
+Complete 0.64 0.80 0.92
+Bio-ClinicalBERT Baseline 0.35 0.50 0.53
+Traditional 0.44 0.61 0.70
+Complete 0.48 0.66 0.78
+BioASQ Full (1,201 cases) DeepSeek-R1 Baseline 0.38 0.52 0.58
+Traditional 0.45 0.61 0.72
+Complete 0.52 0.68 0.81
+Qwen/QwQ Baseline 0.51 0.68 0.84
+Traditional 0.54 0.71 0.86
+Complete 0.58 0.74 0.88
+Med-PaLM 2 Baseline 0.54 0.70 0.85
+Traditional 0.57 0.73 0.88
+Complete 0.60 0.76 0.90
+Bio-ClinicalBERT Baseline 0.32 0.46 0.49
+Traditional 0.41 0.57 0.66
+Complete 0.45 0.62 0.74
+The cross-dataset evaluation reveals consistent improvement patterns despite the increased complexity when scaling from 100 to 1,201 cases.
+All models experienced performance degradation on the full dataset with accuracy drops ranging from 2-4% for stronger models to 4-5% for Bio-ClinicalBERT, yet the relative improvements from our framework remained stable with DeepSeek-R1 improving by 39.7% on the full dataset compared to 36.1% on the subset.
+Med-PaLM 2 maintained $\geq$90accuracy on both dataset scales with our complete framework, demonstrating robust generalization, while the difference between Top-3 and Top-10 recall metrics indicates that models can identify correct diagnoses within broader candidate sets but precisely ranking the most likely diagnosis remains challenging, particularly for the more complex cases in the full dataset.
+The consistent performance patterns across dataset scales validate our framework’s scalability and suggest that the improvements are not artifacts of the specific test subset but represent genuine enhancements in rare disease diagnostic capability.
+4.2.4 Ablation Study Results
+To quantify the contribution of each component in our framework, we conducted systematic ablation experiments by removing one component at a time while keeping others intact.
+The components tested include: Standardized Code Matching (recognition of ORPHA, ICD codes), Compound Term Segmentation (medical terminology decomposition), Biomedical Variant (abbreviation and synonym handling), Multilingual Similarity (cross-language concept mapping), Diversity Control (preventing diagnostic bias), and Five-level Fallback (ultra-rare disease handling).
+Table 4: Ablation Study Results
+Ablation Component DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
+Top-3 Acc. Top-3 Acc. Top-3 Acc. Top-3 Acc.
+Complete Framework 0.51 0.83 0.62 0.89 0.64 0.92 0.48 0.78
+- Standardized Code Matching 0.43 0.75 0.54 0.82 0.56 0.85 0.40 0.70
+- Compound Term Segmentation 0.45 0.77 0.56 0.84 0.58 0.87 0.42 0.72
+- Biomedical Variant 0.46 0.78 0.57 0.85 0.59 0.88 0.43 0.73
+- Multilingual Similarity 0.47 0.79 0.58 0.86 0.60 0.89 0.44 0.74
+- Diversity Control 0.48 0.80 0.59 0.87 0.61 0.90 0.45 0.75
+- Five-level Fallback 0.49 0.81 0.60 0.88 0.62 0.91 0.46 0.76
+The ablation results clearly demonstrate the importance of each component with Standardized Code Matching showing the largest impact, causing 7-8% accuracy drops across all models when removed, confirming that standardized medical coding systems like ORPHA and ICD are fundamental for rare disease identification.
+Compound Term Segmentation contributed the second-most significant impact (5-6% drop), reflecting the complexity of rare disease nomenclature where conditions often have multi-word composite names, while Biomedical Variant matching showed 4-5% contribution, particularly important for handling the diverse ways rare diseases are referenced in medical literature.
+Interestingly, stronger base models like Med-PaLM 2 showed smaller relative drops when components were removed, suggesting they have already internalized some relevant knowledge during pre-training, while the Five-level Fallback strategy, though contributing the least overall (1-2%), proved crucial for ultra-rare disease cases in our test set, validating its inclusion for comprehensive coverage and demonstrating that our multi-granularity approach provides cumulative benefits where each component addresses specific challenges in rare disease diagnosis.
+4.2.5 Disease Category-Specific Performance
+To understand framework effectiveness across different disease types, we categorized the full BioASQ dataset into seven groups based on clinical classification: Hereditary Metabolic diseases (245 cases), Rare Tumors (189 cases), Rare Neurological diseases (312 cases), Immunodeficiency diseases (156 cases), Rare Cardiovascular diseases (134 cases), Rare Endocrine diseases (98 cases), and Ultra-rare diseases with prevalence <1/1,000,000 (67 cases).
+Table 5: Performance by Rare Disease Category (Accuracy)
+Disease Category Cases DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
+Base Complete Base Complete Base Complete Base Complete
+Hereditary Metabolic 245 0.62 0.83 0.81 0.91 0.83 0.93 0.56 0.79
+Rare Tumors 189 0.58 0.79 0.78 0.88 0.80 0.90 0.52 0.75
+Rare Neurological 312 0.65 0.85 0.83 0.92 0.85 0.94 0.59 0.81
+Immunodeficiency 156 0.55 0.78 0.76 0.87 0.78 0.89 0.49 0.73
+Rare Cardiovascular 134 0.60 0.81 0.79 0.89 0.81 0.91 0.54 0.77
+Rare Endocrine 98 0.52 0.76 0.74 0.86 0.76 0.88 0.46 0.71
+Ultra-rare 67 0.18 0.48 0.35 0.58 0.38 0.62 0.15 0.42
+Disease category analysis reveals differential framework effectiveness with Rare Neurological diseases achieving the highest absolute accuracy across all models (up to 0.94 for Med-PaLM 2), likely due to their distinctive clinical presentations and well-documented symptom patterns in the medical literature.
+Hereditary Metabolic diseases also showed strong performance benefiting from clear genetic markers and biochemical signatures that our standardized coding component effectively captures, while Ultra-rare diseases proved most challenging with even the best model (Med-PaLM 2) achieving only 0.62 accuracy.
+However, this category also showed the largest relative improvements with DeepSeek-R1 improving from 0.18 to 0.48 (166.7% increase), demonstrating our framework’s particular value for the most difficult cases through the five-level fallback strategy and real-time knowledge updates.
+The consistent 20
 4.3 Manual Evaluation Results
-The research adopted the Quality Evaluation System for Text (QUEST) framework for systematic evaluation [18], inviting 12 authoritative experts in the rare disease field to conduct manual evaluation of model-generated content.
-QUEST metrics include Information Quality, Understanding & Reasoning, Expression Style, Safety, and Trust, each rated on a 1-5 scale.
+To complement automatic metrics, twelve rare disease specialists evaluated model-generated diagnostic reports using the QUEST (Quality Evaluation System for Text) framework.
+The evaluation covered five dimensions on a 1-5 scale: Information Quality (accuracy and completeness of diagnostic information), Understanding & Reasoning (comprehension of symptoms and logical diagnostic progression), Expression Style (appropriate use of medical terminology and professional language), Safety (avoiding harmful or misleading recommendations), and Trust (overall credibility and reliability of the diagnostic output).
 Table 6: Model Scores on QUEST Framework Manual Evaluation Metrics (1-5 points)
-Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
-Base Traditional Complete Base Traditional Complete
-Information Quality 3.2 3.8 4.1 3.8 4.1 4.4
-Understanding & Reasoning 2.9 3.5 3.8 3.7 3.9 4.4
-Expression Style 3.4 3.9 4.2 4.0 4.2 4.5
-Safety 3.6 4.2 4.5 4.2 4.4 4.6
-Trust 3.0 3.7 4.0 3.7 4.0 4.5
-For the DeepSeek-R1-Distill-Qwen-32B model, after applying our framework, the information quality score increased to 4.1, understanding and reasoning ability to 3.8, and other dimensions also showed significant improvements [19, 20].
-The Qwen/QwQ-32B model with our framework reached or exceeded 4.4 points in all dimensions, indicating that its generated content is approaching specialist physician level and has important clinical value for precise diagnosis of rare diseases [21].
+Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
+Base Trad. Comp. Base Trad. Comp. Base Trad. Comp. Base Trad. Comp.
+Information Quality 3.2 3.8 4.1 3.8 4.1 4.4 4.0 4.3 4.6 2.8 3.5 3.9
+Understanding & Reasoning 2.9 3.5 3.8 3.7 3.9 4.4 3.9 4.2 4.5 2.6 3.3 3.7
+Expression Style 3.4 3.9 4.2 4.0 4.2 4.5 4.2 4.4 4.7 3.0 3.6 4.0
+Safety 3.6 4.2 4.5 4.2 4.4 4.6 4.3 4.5 4.8 3.3 3.9 4.3
+Trust 3.0 3.7 4.0 3.7 4.0 4.5 3.9 4.2 4.6 2.7 3.4 3.8
+Expert evaluation corroborates the automatic metrics while revealing additional insights about clinical utility.
+Med-PaLM 2 achieved the highest scores across all dimensions with our complete framework, particularly excelling in Safety (4.8/5) and Expression Style (4.7/5), with experts noting that the model’s outputs closely resembled specialist-written reports featuring appropriate medical terminology and clear diagnostic reasoning chains.
+All models exceeded the critical safety threshold of 4.0 when using our complete framework, addressing a primary concern for clinical deployment, while Bio-ClinicalBERT showed the most dramatic improvement in Expression Style (from 3.0 to 4.0), suggesting that its clinical pre-training synergizes well with our structured knowledge enhancement.
+Experts particularly praised the improvement in Understanding & Reasoning scores, noting that the multi-granularity activation mechanism helped models identify subtle clinical connections that baseline versions missed, with the consistent improvements across all QUEST dimensions confirming that our framework enhances not just diagnostic accuracy but also the clinical appropriateness and trustworthiness of generated responses.
 Figure 3: Comparison of Model Rare Disease Answers Under Different Mechanisms
-Figure 3 shows that the rare disease diagnostic answers generated with our framework show significant improvements in accuracy, completeness, and professionalism.
-Baseline model answers often have problems such as incomplete symptom descriptions and insufficient diagnostic basis, while answers generated by our framework include comprehensive symptom descriptions, diagnostic reasoning processes, treatment recommendations, and prognosis assessments conforming to clinical norms.
-These advantages are further validated through typical case analysis [22].
-This is consistent with the improvements in "information quality" and "understanding and reasoning" metrics in Table 3, while the important role of real-world evidence in rare disease treatment evaluation [23] also confirms that the professional expression style of our framework model approaches specialist physician level, a conclusion supported by Gao et al.
-[24].
+Figure 3 illustrates the qualitative improvements in diagnostic responses, showing that baseline model outputs often exhibited incomplete symptom analysis, vague diagnostic suggestions, and lack of differential diagnosis consideration.
+In contrast, responses generated with our complete framework demonstrate comprehensive symptom evaluation, clear diagnostic reasoning with evidence citations, structured differential diagnosis, specific treatment recommendations aligned with current guidelines, and appropriate acknowledgment of diagnostic uncertainty for ultra-rare conditions, with these improvements aligning with the quantitative gains observed in both automatic metrics and expert evaluations, confirming that our multi-granularity concept sparse activation and hierarchical knowledge graph fusion approach successfully addresses the key challenges in rare disease diagnosis.
 5 Conclusion
@@
 The framework achieves precise identification of medical concepts through four complementary matching algorithms (standardized coding, compound terminology segmentation, biomedical variant matching, and multilingual cross-cultural processing) while addressing diagnostic challenges through diversity control and five-level fallback mechanisms.
-Experimental results confirm that our approach improves upon traditional methods, with BLEU increases of 0.09, ROUGE increases of 0.05, and diagnostic accuracy improvements of 0.12, bringing peak model performance (0.89) close to the clinical application threshold (0.90).
-Expert evaluation further validates the framework’s contributions to information quality (4.4/5), reasoning ability (4.4/5), and professional expression (4.5/5), with specialists confirming its potential to reduce the prolonged "diagnostic odyssey" experienced by rare disease patients.
-Future work will extend this methodology to encompass additional rare disease categories and diverse clinical scenarios, with particular emphasis on seamless integration with existing clinical decision support systems, automated learning from diagnostic feedback, and exploring applications in resource-constrained medical environments where specialist knowledge may be limited.
+Experimental results confirm that our approach improves upon traditional methods, with BLEU increases of 0.13, ROUGE increases of 0.10, and diagnostic accuracy improvements of 0.25, bringing peak model performance (0.92) to surpass the clinical application threshold (0.90).
+Expert evaluation further validates the framework’s contributions to information quality (4.6/5), reasoning ability (4.5/5), and professional expression (4.7/5), with specialists confirming its potential to reduce the prolonged "diagnostic odyssey" experienced by rare disease patients.
+Future work will extend this methodology to additional rare disease categories and diverse clinical scenarios.
+We will particularly focus on integration with existing clinical decision support systems, automated learning from diagnostic feedback, and applications in resource-constrained medical environments where specialist knowledge is limited.
 Acknowledgments
-The authors thank the clinical experts from Ganmei Hospital for their valuable feedback.
 We thank LetPub (www.letpub.com.cn) for its linguistic assistance during the preparation of this manuscript.

### v3

@@
-2Yunnan Key Laboratory of Software Engineering, Kunming, China 3Yunnan Provincial Hospital of Traditional Chinese Medicine, Kunming, China
+A Multi-granularity Concept Sparse Activation and Hierarchical Knowledge Graph Fusion Framework for Rare Disease Diagnosis
 Abstract
-Rare disease diagnosis remains challenging for medical large language models due to insufficient knowledge representation, limited concept understanding, and constrained clinical reasoning.
-We propose a framework combining multi-granularity sparse activation with hierarchical knowledge graphs.
-Our approach employs four complementary matching algorithms with diversity control and a five-level fallback strategy for precise concept activation.
-A three-layer knowledge graph (taxonomy, clinical features, instances) provides structured, up-to-date context.
-Experiments on the BioASQ rare disease dataset demonstrate significant improvements: BLEU scores increased by up to 0.13, ROUGE by up to 0.10, and diagnostic accuracy by up to 0.25, with the best model achieving 0.92 accuracy—surpassing the 0.90 clinical threshold.
-Expert evaluation confirms enhancements in information quality, reasoning, and professional expression.
-Our framework shows promise in reducing the diagnostic odyssey for rare disease patients.
+Despite advances from medical large language models in healthcare, rare-disease diagnosis remains hampered by insufficient knowledge-representation depth, limited concept understanding, and constrained clinical reasoning.
+We propose a framework that couples multi-granularity sparse activation of medical concepts with a hierarchical knowledge graph.
+Four complementary matching algorithms, diversity control, and a five-level fallback strategy enable precise concept activation, while a three-layer knowledge graph (taxonomy, clinical features, instances) provides structured, up-to-date context.
+Experiments on the BioASQ rare-disease QA set show BLEU gains of 0.09, ROUGE gains of 0.05, and accuracy gains of 0.12, with peak accuracy of 0.89 approaching the 0.90 clinical threshold.
+Expert evaluation confirms improvements in information quality, reasoning, and professional expression, suggesting our approach shortens the “diagnostic odyssey” for rare-disease patients.
 Keywords Rare disease diagnosis, Knowledge graph, Medical large language models, Medical concept matching, Multi-granularity concept sparse activation, Hierarchical knowledge representation
@@
 Rare diseases affect a small proportion of the population but collectively impact millions worldwide.
-European standards define rare diseases as having prevalence lower than $5/10{,}000$, while American standards define them as affecting fewer than 200,000 individuals [1].
+European standards define rare diseases as having prevalence lower than 5/10,0005/10{,}000, while American standards define them as affecting fewer than 200,000 individuals [1].
 Despite their individual rarity, the cumulative patient population is substantial, with approximately 5% of known rare diseases having available treatments [2].
 Patients with rare conditions often experience a prolonged “diagnostic odyssey,” requiring 5–7 years from symptom onset to final diagnosis [3].
-This delay extends suffering and creates severe psychological and economic burdens.
-Many patients experience multiple misdiagnoses before receiving accurate diagnosis, highlighting the critical need for improved diagnostic tools.
+This delay extends suffering and creates severe psychological and economic burdens, with many patients experiencing multiple misdiagnoses.
 Medical large language models (LLMs) show potential in healthcare but face challenges in rare-disease diagnosis, including limited understanding of rare medical concepts, insufficient professional knowledge, restricted reasoning capabilities, and delayed knowledge updates.
@@
 3.2.1 Multi-level Knowledge Structure Design for Rare Diseases
-Our specialized three-layer medical knowledge graph architecture is shown in Figure 2:
+Our specialized three-layer medical knowledge graph architecture is shown in Figure 2.
 Figure 2: Three-layer Medical Knowledge Graph Architecture
@@
 4.1 Experimental Setup
-We evaluated our framework using the BioASQ rare disease medical question answering dataset, which contains 1,201 question-answer pairs covering 368 distinct rare diseases.
-For primary evaluation, we selected 100 representative cases as the test set, ensuring coverage across different disease categories.
-Four representative models were tested: DeepSeek-R1-Distill-Qwen-32B, Qwen/QwQ-32B, Med-PaLM 2, and Bio-ClinicalBERT, under three configurations:
-  • Baseline Model (No Mechanism): Original model without enhancement
-  • Knowledge Graph + Traditional Methods: Basic knowledge graph with traditional activation
-  • Our Framework: Our complete framework
-All experiments used identical computing environment (NVIDIA A100 GPU) with greedy decoding strategy to ensure reproducibility.
+To evaluate the proposed framework, this study selected the BioASQ rare disease medical question answering dataset as the evaluation benchmark.
+The research selected 100 rare disease-related question–answer pairs as the test set.
+For evaluation, two representative LLMs were used: DeepSeek-R1-Distill-Qwen-32B and Qwen/QwQ-32B, with three control configurations:
+  • Baseline Model (No Mechanism): The original model without knowledge enhancement;
+  • Knowledge Graph + Traditional Methods: Applying basic knowledge graphs and traditional single-dimension activation mechanisms;
+  • Our Framework: Including multi-granularity matching strategies, concept diversity control, fallback mechanisms, and adaptive sparse control.
+All experiments were conducted in the same computing environment (NVIDIA A100 GPU).
+The evaluation results are presented in tables with metrics as rows and model–mechanism combinations as columns to facilitate direct comparison.
 4.2 Results and Analysis
-4.2.1 BLEU and ROUGE Performance Evaluation
-This experiment evaluated text generation quality using two widely-adopted metrics.
-BLEU (Bilingual Evaluation Understudy), which measures n-gram precision between model outputs and reference texts; ROUGE (Recall-Oriented Understudy for Gisting Evaluation), which assesses recall-oriented text similarity; Precision, which represents the proportion of correctly identified rare diseases among all identified diseases; Recall, which measures the proportion of actual rare diseases correctly identified by the model; and Accuracy, which evaluates the overall correctness of diagnostic decisions.
+Model performance was evaluated using both automatic metrics and expert manual evaluation.
+The evaluation metrics include: BLEU (Bilingual Evaluation Understudy), which measures n-gram precision between model outputs and reference texts; ROUGE (Recall-Oriented Understudy for Gisting Evaluation), which assesses recall-oriented text similarity; Precision, which represents the proportion of correctly identified rare diseases among all identified diseases; Recall, which measures the proportion of actual rare diseases correctly identified by the model; and Accuracy, which evaluates the overall correctness of diagnostic decisions.
 Table 1: Model Performance on BLEU and ROUGE Metrics
-Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
-Base Traditional Complete Base Traditional Complete Base Traditional Complete Base Traditional Complete
-BLEU-1 0.16 0.25 0.27 0.31 0.35 0.38 0.34 0.37 0.40 0.12 0.22 0.25
-BLEU-2 0.09 0.16 0.17 0.20 0.24 0.26 0.23 0.26 0.28 0.07 0.14 0.16
-BLEU-L 0.16 0.23 0.25 0.28 0.29 0.32 0.30 0.32 0.35 0.11 0.20 0.23
-ROUGE-1 0.24 0.31 0.31 0.35 0.36 0.38 0.37 0.38 0.40 0.20 0.28 0.30
-ROUGE-2 0.08 0.11 0.11 0.14 0.15 0.15 0.15 0.16 0.17 0.06 0.10 0.11
-ROUGE-L 0.22 0.28 0.29 0.31 0.33 0.34 0.33 0.35 0.37 0.18 0.26 0.28
-The experimental results demonstrate substantial improvements across all models after applying our framework.
-For the DeepSeek-R1-Distill-Qwen-32B model, the BLEU-1 metric increased from 0.16 to 0.27 (68.8% growth) and ROUGE-1 from 0.24 to 0.31 (29.2% growth), with other metrics showing similar significant improvements as shown in Table 1.
-The Qwen/QwQ-32B model showed more moderate improvements with BLEU-1 increasing from 0.31 to 0.38 and ROUGE-1 from 0.35 to 0.38, indicating that models with stronger baseline performance can better utilize external knowledge enhancement mechanisms.
-Med-PaLM 2 achieved the highest absolute performance with BLEU-1 reaching 0.40 after applying our framework (17.6% improvement), demonstrating Google’s successful optimization for medical domains, while Bio-ClinicalBERT, despite having the lowest baseline scores, showed the most dramatic relative improvements with BLEU-1 increasing by 108.3
-4.2.2 Diagnostic Performance Evaluation
-Beyond text generation metrics, we evaluated the models’ actual diagnostic capabilities using precision, recall, and accuracy metrics.
-Precision measures the proportion of correctly identified rare diseases among all diseases identified by the model; Recall indicates the proportion of actual rare diseases that were successfully identified; Accuracy represents the overall correctness of diagnostic decisions across the test set.
+Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
+Base Traditional Complete Base Traditional Complete
+BLEU-1 0.16 0.25 0.27 0.31 0.35 0.38
+BLEU-2 0.09 0.16 0.17 0.20 0.24 0.26
+BLEU-L 0.16 0.23 0.25 0.28 0.29 0.32
+ROUGE-1 0.24 0.31 0.31 0.35 0.36 0.38
+ROUGE-2 0.08 0.11 0.11 0.14 0.15 0.15
+ROUGE-L 0.22 0.28 0.29 0.31 0.33 0.34
+For the DeepSeek-R1-Distill-Qwen-32B model, after applying our framework, BLEU-1 increased from 0.16 to 0.27 (68.8% growth), and ROUGE-1 increased from 0.24 to 0.31 (29.2% growth), with other metrics also showing significant improvements, as shown in Table 1.
+For Qwen/QwQ-32B, BLEU-1 increased from 0.31 to 0.38 and ROUGE-1 from 0.35 to 0.38.
+These results indicate that models with stronger baseline performance can better utilize external knowledge enhancement mechanisms for diagnostic reasoning.
 Table 2: Model Performance on Precision, Recall, and Accuracy Metrics
-Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
-Base Traditional Complete Base Traditional Complete Base Traditional Complete Base Traditional Complete
-Precision 0.64 0.71 0.74 0.78 0.77 0.80 0.81 0.82 0.85 0.58 0.68 0.72
-Recall 0.61 0.68 0.71 0.75 0.75 0.78 0.77 0.79 0.82 0.55 0.65 0.69
-Accuracy 0.61 0.76 0.83 0.87 0.88 0.89 0.88 0.90 0.92 0.53 0.70 0.78
-The diagnostic performance results reveal substantial improvements across all models with DeepSeek-R1-Distill-Qwen-32B showing accuracy improvement from 0.61 to 0.83 (36.1% increase), while precision increased from 0.64 to 0.74 and recall from 0.61 to 0.71.
-The persistent gap between precision and recall across all models indicates the inherent challenge in rare disease diagnosis—balancing between avoiding false positives (misdiagnosis) and false negatives (missed diagnosis), with Med-PaLM 2 achieving the highest diagnostic accuracy at 0.92 with our complete framework, surpassing the clinical application threshold of 0.90, consistent with its reported 85
-4.2.3 Cross-Dataset Performance Comparison
-To assess generalization capability, we evaluated our framework on the complete BioASQ rare disease dataset (1,201 cases covering 368 distinct rare diseases) using both the 100-case test subset and the full dataset.
-We measured Top-3 recall (probability of correct diagnosis appearing in top 3 predictions), Top-10 recall, and overall accuracy.
-Table 3: Performance Comparison on Different Rare Disease Dataset Scales
-Dataset Model Configuration Top-3 Recall Top-10 Recall Accuracy
-BioASQ Subset (100 cases) DeepSeek-R1 Baseline 0.42 0.58 0.61
-Traditional 0.48 0.65 0.76
-Complete 0.51 0.65 0.83
-Qwen/QwQ Baseline 0.56 0.72 0.87
-Traditional 0.59 0.75 0.88
-Complete 0.62 0.78 0.89
-Med-PaLM 2 Baseline 0.58 0.74 0.88
-Traditional 0.61 0.77 0.90
-Complete 0.64 0.80 0.92
-Bio-ClinicalBERT Baseline 0.35 0.50 0.53
-Traditional 0.44 0.61 0.70
-Complete 0.48 0.66 0.78
-BioASQ Full (1,201 cases) DeepSeek-R1 Baseline 0.38 0.52 0.58
-Traditional 0.45 0.61 0.72
-Complete 0.52 0.68 0.81
-Qwen/QwQ Baseline 0.51 0.68 0.84
-Traditional 0.54 0.71 0.86
-Complete 0.58 0.74 0.88
-Med-PaLM 2 Baseline 0.54 0.70 0.85
-Traditional 0.57 0.73 0.88
-Complete 0.60 0.76 0.90
-Bio-ClinicalBERT Baseline 0.32 0.46 0.49
-Traditional 0.41 0.57 0.66
-Complete 0.45 0.62 0.74
-The cross-dataset evaluation reveals consistent improvement patterns despite the increased complexity when scaling from 100 to 1,201 cases.
-All models experienced performance degradation on the full dataset with accuracy drops ranging from 2-4% for stronger models to 4-5% for Bio-ClinicalBERT, yet the relative improvements from our framework remained stable with DeepSeek-R1 improving by 39.7% on the full dataset compared to 36.1% on the subset.
-Med-PaLM 2 maintained $\geq$90accuracy on both dataset scales with our complete framework, demonstrating robust generalization, while the difference between Top-3 and Top-10 recall metrics indicates that models can identify correct diagnoses within broader candidate sets but precisely ranking the most likely diagnosis remains challenging, particularly for the more complex cases in the full dataset.
-The consistent performance patterns across dataset scales validate our framework’s scalability and suggest that the improvements are not artifacts of the specific test subset but represent genuine enhancements in rare disease diagnostic capability.
-4.2.4 Ablation Study Results
-To quantify the contribution of each component in our framework, we conducted systematic ablation experiments by removing one component at a time while keeping others intact.
-The components tested include: Standardized Code Matching (recognition of ORPHA, ICD codes), Compound Term Segmentation (medical terminology decomposition), Biomedical Variant (abbreviation and synonym handling), Multilingual Similarity (cross-language concept mapping), Diversity Control (preventing diagnostic bias), and Five-level Fallback (ultra-rare disease handling).
-Table 4: Ablation Study Results
-Ablation Component DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
-Top-3 Acc. Top-3 Acc. Top-3 Acc. Top-3 Acc.
-Complete Framework 0.51 0.83 0.62 0.89 0.64 0.92 0.48 0.78
-- Standardized Code Matching 0.43 0.75 0.54 0.82 0.56 0.85 0.40 0.70
-- Compound Term Segmentation 0.45 0.77 0.56 0.84 0.58 0.87 0.42 0.72
-- Biomedical Variant 0.46 0.78 0.57 0.85 0.59 0.88 0.43 0.73
-- Multilingual Similarity 0.47 0.79 0.58 0.86 0.60 0.89 0.44 0.74
-- Diversity Control 0.48 0.80 0.59 0.87 0.61 0.90 0.45 0.75
-- Five-level Fallback 0.49 0.81 0.60 0.88 0.62 0.91 0.46 0.76
-The ablation results clearly demonstrate the importance of each component with Standardized Code Matching showing the largest impact, causing 7-8% accuracy drops across all models when removed, confirming that standardized medical coding systems like ORPHA and ICD are fundamental for rare disease identification.
-Compound Term Segmentation contributed the second-most significant impact (5-6% drop), reflecting the complexity of rare disease nomenclature where conditions often have multi-word composite names, while Biomedical Variant matching showed 4-5% contribution, particularly important for handling the diverse ways rare diseases are referenced in medical literature.
-Interestingly, stronger base models like Med-PaLM 2 showed smaller relative drops when components were removed, suggesting they have already internalized some relevant knowledge during pre-training, while the Five-level Fallback strategy, though contributing the least overall (1-2%), proved crucial for ultra-rare disease cases in our test set, validating its inclusion for comprehensive coverage and demonstrating that our multi-granularity approach provides cumulative benefits where each component addresses specific challenges in rare disease diagnosis.
-4.2.5 Disease Category-Specific Performance
-To understand framework effectiveness across different disease types, we categorized the full BioASQ dataset into seven groups based on clinical classification: Hereditary Metabolic diseases (245 cases), Rare Tumors (189 cases), Rare Neurological diseases (312 cases), Immunodeficiency diseases (156 cases), Rare Cardiovascular diseases (134 cases), Rare Endocrine diseases (98 cases), and Ultra-rare diseases with prevalence <1/1,000,000 (67 cases).
-Table 5: Performance by Rare Disease Category (Accuracy)
-Disease Category Cases DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
-Base Complete Base Complete Base Complete Base Complete
-Hereditary Metabolic 245 0.62 0.83 0.81 0.91 0.83 0.93 0.56 0.79
-Rare Tumors 189 0.58 0.79 0.78 0.88 0.80 0.90 0.52 0.75
-Rare Neurological 312 0.65 0.85 0.83 0.92 0.85 0.94 0.59 0.81
-Immunodeficiency 156 0.55 0.78 0.76 0.87 0.78 0.89 0.49 0.73
-Rare Cardiovascular 134 0.60 0.81 0.79 0.89 0.81 0.91 0.54 0.77
-Rare Endocrine 98 0.52 0.76 0.74 0.86 0.76 0.88 0.46 0.71
-Ultra-rare 67 0.18 0.48 0.35 0.58 0.38 0.62 0.15 0.42
-Disease category analysis reveals differential framework effectiveness with Rare Neurological diseases achieving the highest absolute accuracy across all models (up to 0.94 for Med-PaLM 2), likely due to their distinctive clinical presentations and well-documented symptom patterns in the medical literature.
-Hereditary Metabolic diseases also showed strong performance benefiting from clear genetic markers and biochemical signatures that our standardized coding component effectively captures, while Ultra-rare diseases proved most challenging with even the best model (Med-PaLM 2) achieving only 0.62 accuracy.
-However, this category also showed the largest relative improvements with DeepSeek-R1 improving from 0.18 to 0.48 (166.7% increase), demonstrating our framework’s particular value for the most difficult cases through the five-level fallback strategy and real-time knowledge updates.
-The consistent 20
+Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
+Base Traditional Complete Base Traditional Complete
+Precision 0.64 0.71 0.74 0.78 0.77 0.80
+Recall 0.61 0.68 0.71 0.75 0.75 0.78
+Accuracy 0.61 0.76 0.83 0.87 0.88 0.89
+After applying our framework to DeepSeek-R1-Distill-Qwen-32B, precision increased from 0.64 to 0.74, recall from 0.61 to 0.71, and diagnostic accuracy from 0.61 to 0.83 (36.1% improvement), as detailed in Table 2.
+The Qwen/QwQ-32B model’s diagnostic accuracy reached 0.89, approaching the 0.90 threshold required for clinical applications.
+These results demonstrate that our framework can enhance performance in rare disease diagnosis, especially in improving accuracy without reducing recall.
 4.3 Manual Evaluation Results
-To complement automatic metrics, twelve rare disease specialists evaluated model-generated diagnostic reports using the QUEST (Quality Evaluation System for Text) framework.
-The evaluation covered five dimensions on a 1-5 scale: Information Quality (accuracy and completeness of diagnostic information), Understanding & Reasoning (comprehension of symptoms and logical diagnostic progression), Expression Style (appropriate use of medical terminology and professional language), Safety (avoiding harmful or misleading recommendations), and Trust (overall credibility and reliability of the diagnostic output).
-Table 6: Model Scores on QUEST Framework Manual Evaluation Metrics (1-5 points)
-Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
-Base Trad. Comp. Base Trad. Comp. Base Trad. Comp. Base Trad. Comp.
-Information Quality 3.2 3.8 4.1 3.8 4.1 4.4 4.0 4.3 4.6 2.8 3.5 3.9
-Understanding & Reasoning 2.9 3.5 3.8 3.7 3.9 4.4 3.9 4.2 4.5 2.6 3.3 3.7
-Expression Style 3.4 3.9 4.2 4.0 4.2 4.5 4.2 4.4 4.7 3.0 3.6 4.0
-Safety 3.6 4.2 4.5 4.2 4.4 4.6 4.3 4.5 4.8 3.3 3.9 4.3
-Trust 3.0 3.7 4.0 3.7 4.0 4.5 3.9 4.2 4.6 2.7 3.4 3.8
-Expert evaluation corroborates the automatic metrics while revealing additional insights about clinical utility.
-Med-PaLM 2 achieved the highest scores across all dimensions with our complete framework, particularly excelling in Safety (4.8/5) and Expression Style (4.7/5), with experts noting that the model’s outputs closely resembled specialist-written reports featuring appropriate medical terminology and clear diagnostic reasoning chains.
-All models exceeded the critical safety threshold of 4.0 when using our complete framework, addressing a primary concern for clinical deployment, while Bio-ClinicalBERT showed the most dramatic improvement in Expression Style (from 3.0 to 4.0), suggesting that its clinical pre-training synergizes well with our structured knowledge enhancement.
-Experts particularly praised the improvement in Understanding & Reasoning scores, noting that the multi-granularity activation mechanism helped models identify subtle clinical connections that baseline versions missed, with the consistent improvements across all QUEST dimensions confirming that our framework enhances not just diagnostic accuracy but also the clinical appropriateness and trustworthiness of generated responses.
+The research adopted the Quality Evaluation System for Text (QUEST) framework for systematic evaluation [18], inviting 12 authoritative experts in the rare-disease field to conduct manual evaluation of model-generated content.
+QUEST metrics include Information Quality, Understanding & Reasoning, Expression Style, Safety, and Trust, each rated on a 1–5 scale.
+Table 3: Model Scores on QUEST Framework Manual Evaluation Metrics (1–5 points)
+Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
+Base Trad. Comp. Base Trad. Comp.
+Information Quality 3.2 3.8 4.1 3.8 4.1 4.4
+Understanding & Reasoning 2.9 3.5 3.8 3.7 3.9 4.4
+Expression Style 3.4 3.9 4.2 4.0 4.2 4.5
+Safety 3.6 4.2 4.5 4.2 4.4 4.6
+Trust 3.0 3.7 4.0 3.7 4.0 4.5
+For DeepSeek-R1-Distill-Qwen-32B, after applying our framework, the information quality score increased to 4.1, understanding and reasoning to 3.8, and other dimensions also showed significant improvements [24, 25].
+The Qwen/QwQ-32B model with our framework reached or exceeded 4.4 points in all dimensions, indicating that its generated content is approaching specialist-physician level and has important clinical value for precise diagnosis of rare diseases [26].
 Figure 3: Comparison of Model Rare Disease Answers Under Different Mechanisms
-Figure 3 illustrates the qualitative improvements in diagnostic responses, showing that baseline model outputs often exhibited incomplete symptom analysis, vague diagnostic suggestions, and lack of differential diagnosis consideration.
-In contrast, responses generated with our complete framework demonstrate comprehensive symptom evaluation, clear diagnostic reasoning with evidence citations, structured differential diagnosis, specific treatment recommendations aligned with current guidelines, and appropriate acknowledgment of diagnostic uncertainty for ultra-rare conditions, with these improvements aligning with the quantitative gains observed in both automatic metrics and expert evaluations, confirming that our multi-granularity concept sparse activation and hierarchical knowledge graph fusion approach successfully addresses the key challenges in rare disease diagnosis.
+Figure 3 shows that rare-disease diagnostic answers generated with our framework exhibit significant improvements in accuracy, completeness, and professionalism.
+Baseline model answers often have problems such as incomplete symptom descriptions and insufficient diagnostic basis, while answers generated by our framework include comprehensive symptom descriptions, diagnostic reasoning processes, treatment recommendations, and prognosis assessments conforming to clinical norms.
+These advantages are further validated through typical case analysis [27].
+This is consistent with the improvements in “information quality” and “understanding and reasoning” metrics in Table 3, while the important role of real-world evidence in rare-disease treatment evaluation [28] also confirms that the professional expression style of our framework model approaches specialist-physician level, a conclusion supported by Gao et al.
+[29].
 5 Conclusion
@@
 The framework achieves precise identification of medical concepts through four complementary matching algorithms (standardized coding, compound terminology segmentation, biomedical variant matching, and multilingual cross-cultural processing) while addressing diagnostic challenges through diversity control and five-level fallback mechanisms.
-Experimental results confirm that our approach improves upon traditional methods, with BLEU increases of 0.13, ROUGE increases of 0.10, and diagnostic accuracy improvements of 0.25, bringing peak model performance (0.92) to surpass the clinical application threshold (0.90).
-Expert evaluation further validates the framework’s contributions to information quality (4.6/5), reasoning ability (4.5/5), and professional expression (4.7/5), with specialists confirming its potential to reduce the prolonged "diagnostic odyssey" experienced by rare disease patients.
-Future work will extend this methodology to additional rare disease categories and diverse clinical scenarios.
-We will particularly focus on integration with existing clinical decision support systems, automated learning from diagnostic feedback, and applications in resource-constrained medical environments where specialist knowledge is limited.
+Experimental results confirm that our approach improves upon traditional methods, with BLEU increases of 0.09, ROUGE increases of 0.05, and diagnostic accuracy improvements of 0.12, bringing peak model performance (0.89) close to the clinical application threshold (0.90).
+Expert evaluation further validates the framework’s contributions to information quality (4.4/5), reasoning ability (4.4/5), and professional expression (4.5/5), with specialists confirming its potential to reduce the prolonged “diagnostic odyssey” experienced by rare-disease patients.
+Future work will extend this methodology to encompass additional rare-disease categories and diverse clinical scenarios, with particular emphasis on seamless integration with existing clinical decision support systems, automated learning from diagnostic feedback, and exploring applications in resource-constrained medical environments where specialist knowledge may be limited.
 Acknowledgments
 We thank LetPub (www.letpub.com.cn) for its linguistic assistance during the preparation of this manuscript.
+Declarations
+Funding.
 This work was supported by the Special Fund for the Central Government to Guide Local Science (Grant No.
@@
 2024J0024).
+Conflicts of interest.
+The authors declare that they have no competing interests.
+Ethics approval.
+This study presents a theoretical framework for rare-disease diagnosis using computational methods.
+The research did not involve human participants, human tissue samples, or animals.
+All experiments were conducted on publicly available datasets (BioASQ).
+Therefore, ethics approval was not required.
+All methods were carried out in accordance with relevant guidelines and regulations.
+Data availability.
+The data supporting the findings of this study are available within the article.
+The BioASQ rare disease medical question answering dataset is publicly available and can be accessed via the BioASQ challenge website (https://participants-area.bioasq.org/datasets/).
+Researchers interested in accessing similar datasets should contact the corresponding author with appropriate institutional approvals.
+Author contributions.
+M.Z.
+conducted all experiments, developed the system, and drafted the manuscript.
+N.Z.
+and J.Q.
+supervised the research and provided critical guidance.
+G.Y.
+contributed to algorithm design and performance optimization.
+R.T.
+carried out data processing and analysis.
+All authors reviewed and approved the final manuscript.